OpenStudy (ianchondo):

Find the vertical asymptotes (if any) of the graph of the function. (Use n as an arbitrary integer if necessary. If an answer does not exist, enter DNE.) s(t) = 2t/sin(t)

zepdrix (zepdrix):

Hey Ian! Welcome to OpenStudy!\[\large\rm s(t)=\frac{2t}{\sin t}\]We have asymptotic behavior when the bottom of this fraction is near zero.

zepdrix (zepdrix):

\[\large\rm \sin t=0\]Do you know how to solve for the values of t from this point?

OpenStudy (ianchondo):

Nope, could you please explain. Thank you.

OpenStudy (ianchondo):

the answer that the book explains is like n times pi, I thought that mean 2 times pi.

zepdrix (zepdrix):

Do you remember anything about your unit circle? sine represents your y-coordinate.

zepdrix (zepdrix):

these are the two locations (within one rotation) where we have no y-coordinate.